Progress Toward the Milestones of the enclosed schedule:
NOTE: Meetings are held every Tuesday in CEBAF Center,
Room #A110, at 11 a.m., unless otherwise notified.
NOTE: A reminder to all principals to submit a brief weekly progress to D. Gilchrist each Friday by noon
ATTENDEES: S. Benson, C. Bohn, L. Dillon-Townes, D. Engwall,
J. Fugitt, K. Jordan, G. Krafft, D. Kehne, K. Hovater,
W. Merz, G. Neil, C. Piller, K. Rode, M. Shinn, T. Siggins,
The wave guide to pipe 50 kW power to the Cryomodule Test Cave
is almost ready. They want to run next month. K. Jordan doesn't
want to release power from our system (as controlled by our safety
system) until the Test Cave is brought up to present standards
of redundancy. This will have to be resolved with C. Reese.
Safety Action Items:
New Safety Action Items:
Last week, two staff members were able to get into the laser
enclosure without goggles while the laser was on because one of
them knew the entry code. Neither were on the qualified personnel
list. Both ignored the lighted warning sign in the entrance alcove.
No injuries resulted. P. Hunt is in charge of the investigation.
In response, the procedures are being tightened: 1) No one gets
the entry code without being on the qualified personnel list;
2) The code will be changed with every 6 month recertification;
3) The List of Qualified Personnel is posted at the entry keypad
and 4) the lighted warning sign in the entrance alcove is enhanced
with a yellow strobe light .
Completed Safety Action Items:
K. Jordan - ITS test cave, Gas lines and taps on back wall not
Determine the need for additional personnel required to match or beat the Schedule. K. Jordan is still not getting the 1 A trim cards he needed for magnet powering. They are promised for tomorrow.
T. Siggins - Write the tricks of polishing learned at SLAC. No action.
G. Biallas - Get cleaning started in the area on a regular basis. G. Biallas will set up a roster for cleaning in the future. No action.
G. Biallas - Define Nomenclature document Definitions have to get to T. Hassler. No action.
J. Susta/M. Wiseman - Alter the permanent LCW system for the Klystrons such that the pressure remains low and put a relief in the waveguides in case a circulator window breaks. System for the second Klystron is being plumbed. The waveguide relief is needed this week if it could be assembled.
G. Biallas - Finish the LCW System: Flow & Temp reading, Level indicators, UV Lamp - UV lamp needs to be commissioned, full qualification walk through needed
D. Kehne - Have simulations or bracketed simulations of the 350 keV projected initial running case for the Gun. D. Kehne's is in process.
K. Jordan/S. Benson - Failure analysis for the 10 MeV beam handling. No action
M. Shinn - Put phase noise documentation on the web. In process.
M. Shinn - Get HeNe at low power onto cathode. In process.
All - general cleanup after bake. Checkout systems, straighten out cabling. Partially done.
R. Walker - send out MO for re calibration and provide replacement. In process.
K. Jordan - move Digitels into control room. K. Jordan will supply roll-around, short cabinet for the enclosure.
L. Dillon-Townes - design transition radiation mirror for dump. No progress.
R. Walker - get Parkinson to hook up water, etc. for chopper. R. Walker to see F. Humphry.
D. Kehne - Publish the straw man for the 10 MeV Experiment .
New Action Items:
G. Biallas - integrate the new experimental schedule developed
by Bob Legg into the main schedule.
Completed Action Items:
R. Walker - Arc detector - Complete, he will get it into the wave guide with M. Augustine.
K. Hovater: measure transfer function of the divide by 40 feedback system. Complete.
K. Hovater - test power capacity of feedthru for chopper. Complete,
new probe being mounted with water cooling loop.
Individual Reports of 7/19/96:
The "leak valve" that controls the amount of NF3 used to activate the cathode, that was closed during the bake, leaked atmosphere into the gun volume when the valve's sealed line was opened. The leak to 10 Torr or so destroyed the baked quality of the gun interior and another bake (as well as a new GA disk) is necessary. We decided it would be prudent to replace the six-way cross at the slit position in the beam line that had been sealed with Vacseal and then bake both the line and the gun at the same time. At weeks end, the cross was replaced and realigned, but a new leak was found in the five-way cross in the spectrometer line.
D. Bullard finished the ramp over the waveguides on the mezzanine.
The material for the mid voltage ring for the next stack of ceramics has huge grain size and is unacceptably soft (64 Rb instead of 80Rb). The vendor is making us new rings.
L. Mandelcorn, the consultant from Northrop Grumman concluded that leaks in the ceramics are caused by the physical shock, (propagating to the outside) from the discharge of charge built up within the inner layer of ceramic to the resistive coating. The charge build-up is caused by 300 to 400 kV electrons that are field emitted from the stalk. He suggests several solutions: (1) Replace the non-lossy 0.5% glass alumina with an alumina that is lossy such as the 3% glass alumina. The charge may bleed off within this ceramic and discharges may never happen. C. Sinclair's 400 kV gun with such ceramics had no problems. This is what we are doing already. (2) Coat the stem and ball with coating that reduces field emission by several orders of magnitude. L. Phillips knows a vendor and I am getting a new batch of material to make a new stem and ball. (3) Replace the alumina in the ceramics with materials such as zircon or silica glass that have much lower bulk resistivities that may not even need a resistive coating and that bleed the charge off easily. (4) Insert a shield of metal to intercept the field emitted electrons. (5) Insert a lossy ceramic like mica or silica glass as a shield as in 4 above.
500 keV experiment :
-Aperture has been put into the brazing que.
10 MeV experiment:
- C-Tek shielded bellows configuration will be used.
- Fast valve tested for operation and vacuum. Both tests were successful.
- Reflective mirror defined for the drift tube diagnostics.
- Assembly drawings being processed.
- Most cooling water components have arrived.
Second circulator still has problem, the cooling plate is being built from scratch the second time. the first rebuild failed again because of bad braze. They are going to use a different brazing vendor. Expect to ship in about 10 days.
Continue work on installation of second HPA unit (called # 1)
We plan to test the HOM Filter with full power when we are ready
to resume high power testing.
Worked on interfacing the HP8561E spectrum analyzer to a computer, in order to get high quality plots for a future publication. Had numerous problems, which were traced to the GPIB interface board. HP (the vendor) replaced the board, and I am having more success.
Assisted C. Hovater with transfer function measurements. The width of the resonance is consistent with specifications by Coherent. We will be making further tests next week.
Determined the Opto-Electronics Inc. photodiode was damaged
while transfer measurements were made. It will be replaced, and
we will purchase another fast photodiode system as well.
A new drive probe was built for the Chopper cavity this week with a nickel center conductor. It will be tested early next week.
The additional waveguide elbows are still not finished by the Machine Shop. The elbows are for single and dual Arc Detector Sensor Blocks. Additional 5 mm Lens Holders will have to be remade as the Machine Shop used the 6 mm drawing.
The prototype Arc Detector was assembled, tested, and installed this week. The final units will be assembled and tested when the front and rear panels are available in a few weeks.
A meeting will be held next week to discuss the software for the IR FEL unit.
The HP8644A which is the FEL Master Oscillator, was sent to HP for testing and calibration. It is due back in 4 to 5 weeks.
The Buncher cavity probes were installed. It will be leak checked
and pumped down next week. Manuals for the water skid controllers
ITS Schedule for 7/22/96